OpenStudy (anonymous):

Find f(-2) if f(x)=x⁴+2x²-1

OpenStudy (ash2326):

f(-2)?? substitute x=-2 in f(x) f(-2)= (-2)^4+2(-2)^2-1 =16+2*4-1 =16+8-1 =23